OpenStudy (anonymous):

Write \[3\frac{ 2 }{ 5 }\] in radical form Please help!

OpenStudy (bahrom7893):

\[\sqrt{3^2*\frac{2^2}{5^2}}\]

OpenStudy (bahrom7893):

Just square everything, and then put it all under a square root.

OpenStudy (e.mccormick):

Do you mean \(3^{\frac{2}{5}}\)?

OpenStudy (anonymous):

yes, in radical form

OpenStudy (bahrom7893):

Wait is \[\frac{2}{5}\] in the power?

OpenStudy (e.mccormick):

The top is the power. The bottom is the radical.

OpenStudy (anonymous):

Yes bahrom

OpenStudy (anonymous):

that is the power

OpenStudy (bahrom7893):

Oh, then e.mccormick is right.

OpenStudy (anonymous):

thankyou bahrom ill post more questions

OpenStudy (e.mccormick):